OpenStudy (anonymous):

Nicole has a clay vase shaped as a triangular prism with a base area of 136.9 square inches. It is one third filled with water. If there are 575 cubic inches of water in the vase, what is the height of the vase, rounded to the nearest tenth?

OpenStudy (anonymous):

(1/3) b * h = 575 => h = 3* 575 / 136.9 = 12.6 in